OpenStudy (itz_sid):

For what values of x does the graph of f(x) = x − 2 sin x have a horizontal tangent? (Use n as your integer variable. Enter your answers as a comma-separated list.)

zepdrix (zepdrix):

Hey Sid! Welcome to OpenStudy! Recall that your derivative function will give you the slope of the line tangent to the curve at a given point. So we're concerned with zero slope, horizontal tangent lines. This happens when \(\large\rm f'(x)=0\). Do you understand how to find your derivative function from what was given?

OpenStudy (itz_sid):

Yea so, f'(x) = 1-2Cosx.

OpenStudy (itz_sid):

And then I have to solve for x?

zepdrix (zepdrix):

Ok great. Then set your derivative function equal to 0, this is what it means to look for horizontal tangents,\[\large\rm 0=1-2\cos x\]And then yes, solve for x. Remember some of your trig stuff? :)

OpenStudy (itz_sid):

Yea, haha. so \[\frac{ \pi }{ 2 }\] ?

OpenStudy (itz_sid):

oh wait sorry

OpenStudy (itz_sid):

\[\frac{ \pi }{ 3 }\]

zepdrix (zepdrix):

So umm you ended up with 1/2 on the other side, right? 1/2 is the shorter length... and cosine is the x coordinate... sooo ya ya ya that happens up at pi/3. But within one rotation of the unit circle, there is one other place.

OpenStudy (itz_sid):

Oh yea, and \[\frac{ 5\pi }{ 3 }\]

zepdrix (zepdrix):

Ok, good. We have something else to worry about as well. Our function is not restricted to a single rotation of the unit circle [0,2pi]. We can spin around the circle any number of times and get this same 1/2 value. Example, from pi/3 if we spin 2pi, we land at the same location, giving us the same 1/2 from our cosine, ya?\[\large\rm \cos\left(\frac{\pi}{3}\right)=\cos\left(\frac{\pi}{3}+2\pi\right)=\frac12\]

OpenStudy (itz_sid):

Oh yea. :/

zepdrix (zepdrix):

So we have to allow for full rotations. Any number of full rotations can be added to our solutions.\[\large\rm x=\frac{\pi}{3}+2n \pi\]Where n is an integer value (as the instructions told us to use). That would be part of our solution, ya?

OpenStudy (itz_sid):

Oh i get it. So would that be the answer?

zepdrix (zepdrix):

That would be your first "set of solutions", we have to include the other set as well. The 5pi/3 with rotations allowed.

OpenStudy (itz_sid):

Oh I see

zepdrix (zepdrix):

Just so it's clear, this is one set of solutions, \(\large\rm \dfrac{\pi}{3}+2n\pi\). You would put a comma before writing the other set of solutions.